OpenStudy (anonymous):

Solve 4/5 x- 3 3/4 x= 2 1/2?

mathslover (mathslover):

\(\cfrac{4}{5} x - 3\cfrac{3}{4} x = 2\cfrac{1}{2}\) First of all solve the mixed fractions.

mathslover (mathslover):

We have 2 mix fractions here : \(3\cfrac{3}{4}\) and \(2\cfrac{1}{2}\) : I will solve one for you, @mynamebejew you will have to solve the second mix fraction on your own, ok? \(3\cfrac{3}{4} = \cfrac{3\times 4 + 3}{4} = \cfrac{12+3}{4} = \cfrac{15}{4}\)

mathslover (mathslover):

Now solve \(2\cfrac{1}{2}\) @mynamebejew , can you do that?

OpenStudy (anonymous):

Sorry I went away from keyboard! Would it be 4/2?